# Determination of Magnetic Symmetries by Convergent Beam Electron Diffraction

### v2

@@ I Introduction
Magnetic point and space groups [1], also referred to as Shubnikov groups, provide a fundamental classification underlying all magnetic or spin-related properties of crystalline solids [2], such as ferro-, ferri- and antiferromagnetism [3], altermagnetism [4], topological insulation [5], and semimetallicity [6, 7]. Magnetic space groups extend the crystallographic space groups by combining spatial symmetries with time-reversal symmetry, which connects opposite spin orientations, and provide compatibility conditions for magnetic properties following Neumann’s principle. Therefore, experimental probes of magnetic point-group and space-group symmetries, analogous to probes of crystallographic symmetries, are indispensable for studies of solid-state magnetism, with neutron diffraction being the main probe [8]. A direct determination of the magnetic symmetry from the symmetries of a neutron diffraction pattern is, however, hampered by Friedel’s law that holds for scattering within the Born approximation: Opposite reflections in a neutron diffraction pattern are always symmetric even for noncentrosymmetric space groups. Motivated by this and other limits of neutron diffraction, such as limited spatial resolution and relatively large required crystal sizes, we establish an alternative technique based on electron diffraction in this paper. We introduce a method for the determination of magnetic point groups that allows direct measurement of symmetries with a spatial resolution reaching several nanometers, referred to as “magnetic convergent-beam electron diffraction.”
@@
While the conventional CBED technique focuses solely on structural symmetries of the crystal imprinted on the electrostatic scattering potential, a growing number of electron-scattering studies highlight the impact of the magnetic field. In an early study, Shen and Laughlin [21] showed the breaking of structural CBED symmetries by the presence of the strong uniaxial macroscopic magnetic field in ferromagnetic PrCo5. Later, Loudon [22] revealed the existence of purely magnetic electron-diffraction spots in parallel-beam electron diffraction on antiferromagnetic NiO, where the antiferromagnetic unit cell is doubled compared to the structural one. This work paved the way for the detection of antiferromagnetic order in the Transmission Electron Microscope (TEM). Recently, Kohno et al. [23] demonstrated the sensitivity of atomic-resolution differential phase contrast (another convergent-beam electron-diffraction technique, where first moments of diffraction patterns are analyzed) to the antiferromagnetic structure of Fe2O3. These and similar observations [24, 25, 26] suggest the experimental possibility to detect the generally weak electron-diffraction effects due to atomic magnetic fields [27] in antiferromagnets, which possess no macroscopic magnetic field. The latter is the main impetus for the following generalization of the conventional CBED symmetry analysis by taking into account symmetries of the magnetic crystal structure.
@@ II Electron diffraction at electric and magnetic potentials
Electron beams accelerated by voltages in the $60\,\mathrm{kV}$ to $300\,\mathrm{kV}$ range attain relativistic velocities (e.g., $v\approx 0.8\,c$ at $300\,\mathrm{kV}$ acceleration voltage) and are predominantly scattered by very small angles (less than $100\,\mathrm{mrad}$) relative to the optical axis in a transmission electron microscope. In this small-angle, or paraxial scattering, limit, the effects of spin through spin-orbit coupling and Zeeman coupling are negligible, meaning the spinor nature of the electron wave function can be ignored. As a result, spin does not influence the diffraction, allowing us to treat the electron’s wave function as a scalar. Moreover, the resulting scalar wave function $\psi$ may be decomposed into a fast carrier wave function and an envelope wave function $\Psi$ that is slowly varying along the optical axis, i.e., $\psi(\boldsymbol{r}_{\bot},z)=e^{ik_{0}z}\,\Psi(\boldsymbol{r}_{\bot},z)$ with $\boldsymbol{r}_{\bot}=\left(x,y\right)$ denoting the in-plane slab coordinates and $k_{0}$ the wave number of the fast carrier wave. Here, the time-dependent oscillation has already been separated, i.e., we only consider eigensolutions with a certain electron energy $E$. Inserting the above decomposition into the Klein-Gordon equation for scalar relativistic particles and neglecting a couple of terms that are small due to the high beam energy, the evolution of the envelope wave function along the optical axis can be described by the paraxial Schrödinger equation [28, 29] (see Appendix A for a detailed derivation)
@@
The symmetries of the solution of the above equation in the absence of the vector potential, $\boldsymbol{A}=0$, have been exploited for structural symmetry determination using CBED based on the following argument. A CBED pattern corresponds to the absolute square of the electron wave function in the far field behind the object (coordinates $\boldsymbol{k}_{\perp}=\left(k_{x},k_{y}\right)^{\mathrm{T}}$). Therefore, the CBED discs appear at integer multiples of unit vectors of the crystal’s reciprocal lattice according to Bragg’s law, see Fig. 1. The electron wave function, on the other hand, inherits symmetries from the paraxial Hamiltonian $\hat{H}_{\bot}$ and hence symmetries of the electric scattering potential $\Phi$, which leads to a direct relationship between symmetries of the latter and those of the CBED pattern [30]. Given the similar mathematical structure of the above equation with and without magnetic vector potential, this reasoning may be also adopted to derive the relationship between CBED pattern symmetries and magnetic point-group symmetries. Because of the additional sign change of the magnetic vector potential upon time-reversal, the latter has to be included in the symmetry considerations, when considering CBED on magnetic materials.
The structure of Eq. (1) corresponds to a two-dimensional time-dependent Schrödinger equation, i.e., a $\left(2+1\right)$-dimensional theory, with $z$ and the paraxial Hamiltonian $\hat{H}_{\bot}$ taking the place of the time coordinate and the usual Hamiltonian, respectively [31, 32]. In the following, we will take advantage of this analogy, allowing for a substantial portion of the time-dependent Schrödinger equation’s properties and solutions to be transferred to paraxial scattering.
@@
[…]
@@
In full analogy to time evolution, the $z$ evolution is unitary, conserving the norm of the wave function along $z$, because the Hamiltonian $H_{\bot}$ is hermitian. Moreover, any in-plane symmetries (rotations, mirror reflections, and translations in the $xy$ plane) leave the propagator invariant because the corresponding unitary symmetry transformations $\hat{T}_{\bot}$ commute with the Hamiltonian $\hat{H}_{\bot}$ and hence with the propagator, $[\hat{T}_{\bot},\hat{K}]=0$. Translational symmetries correspond to phase factors in reciprocal space, which are not directly visible as diffraction symmetries but instead lead to absences of diffraction peaks. Hence, only in-plane rotation and mirror symmetries, i.e., the in-plane point-group symmetry of the slab, translate to symmetries of the CBED pattern, provided that the initial probe wave function also has these symmetries. Rotational and mirror symmetries are present for a circularly symmetric converged electron beam used in CBED experiments. To match these beam symmetries with the point group symmetries of the sample (i.e., to align the beam propagation axis with the high-symmetry crystallographic axis of the sample), the beam direction has to be carefully adjusted in the experiment. Here, the beam position (center of mass of the beam) within a unit cell, i.e., whether it strictly coincides with a high-symmetry axis or not, is typically irrelevant if the size of the convergent illumination spans several unit cells.
@@ III Magnetic CBED groups
Starting from the magnetic point group of the bulk crystal, we first need to construct the subgroup of elements that leave the slab geometry invariant, as discussed in Sec. II. This leads to the $10$ two-dimensional structural point groups, in Schoenflies notation, $C_{n}$ and $D_{n}$ with $n=1,2,3,4,6$. It is easy to check explicitly that all elements of these groups commute with z reversal. The fact that z reversal is antiunitary does not affect the commutation relations  [^1] One can find a purely real representation of all group elements and thereby show that the additional complex conjugation introduced by making z reversal antiunitary does not change the commutation relations.. Time reversal commutes with all possible structural transformations in any case [34]. These arguments justify assumptions (c) and (d).
Including the antiunitary time-reversal operation $\mathcal{T}$ leads to $31$ two-dimensional magnetic point groups in total [35]: First, the $10$ groups, $G$, noted above, which do not contain time reversal and are called “monochromatic.” Second, $10$ groups $G+G\mathcal{T}$, which result from adding time reversal $\mathcal{T}$ as a generator. These groups are called “gray.” Third, $11$ groups that are created by identifying a halving subgroup $H$ of a monochromatic group $G$ and forming the group $H+(G\setminus H)\mathcal{T}$, i.e., by multiplying the coset $G\setminus H$ by time reversal. These groups are called “dichromatic.” For an introduction to magnetic point groups see, for example, [36, 37].
@@
double dichromatic with $c_{n}\mathcal{T}$ $Q+Qc_{n}\mathcal{T}+Qm_{x}\Omega+Qc_{n}m_{x}\mathcal{Z}$ $\mathcal{T}$, $\mathcal{Z}$ dichromatic $3$
double dichromatic with $c_{n}\mathcal{Z}$ $Q+Qc_{n}\mathcal{Z}+Qm_{x}\Omega+Qc_{n}m_{x}\mathcal{T}$ $\mathcal{T}$, $\mathcal{Z}$ dichromatic $3$
double dichromatic with $c_{n}\Omega$ $Q+Qc_{n}\Omega+Qm_{x}\mathcal{T}+Qc_{n}m_{x}\mathcal{Z}$ $\mathcal{T}$, $\mathcal{Z}$ dichromatic $3$
@@ IV Relation between magnetic CBED groups and magnetic point groups
We first consider a slab with surface normal oriented along the $[001]$ crystallographic axis, i.e., the fourfold symmetry axis. Thus, this slab orientation is compatible with the point group $4^{\prime}/m$. The mirror reflection in the horizontal plane is then mapped to the $z$-reversal symmetry, which we denote by $1_{R}$, following the conventions from CBED literature [12]. The corresponding magnetic CBED group is thus $4^{\prime}1_{R}$.
On the other hand, a slab with normal along $[2\bar{1}0]$ is incompatible with all operations that involve rotations by $90^{\circ}$. This reduces the symmetry of the slab to the point group $2/m$ (see Fig. 2). Describing this point group within the coordinate system defined by the slab / CBED experiment (i.e., $z$-axis parallel to beam slab normal / beam direction), the mirror symmetry along $c$ becomes a mirror along $y$ and the twofold rotation around $c$ becomes a $z$-reversal combined with a $x$-mirror. Hence, the twofold rotation is mapped to $m_{R}$. Noting that the product of $m$ and $m_{R}$ is a two-fold rotation within the plane multiplied by $z$-reversal, i.e. $2_{R}$, the magnetic CBED group of the $[2\bar{1}0]$ slab turns out to be $2_{R}mm_{R}$.
These examples illustrate the general procedure we apply to generate the complete mapping of all magnetic point groups: First, we generate a representative magnetic structure for a certain magnetic point group. In the second step, we analyze the preserved symmetry operations for a certain slab orientation. Finally, we map the reversal of the coordinate normal to the slab to the antiunitary $z$-reversal. To cover all possibilities, i.e., all symmetry classes of slab orientations, we probe all classes of orientations that yield different conventional CBED symmetries, as tabulated in the literature [12]. It turns out that all possible 125 magnetic CBED groups from Sec. III are generated by applying the above algorithm for all 122 three-dimensional magnetic point groups. The above procedure was carried out with the help of a software package provided by De Graef [38].
@@ V Computational examples
High-resolution TEM data as well as electron-diffraction data from magnetic crystal structures can be reliably and quantitatively reproduced by numerical integration of Eq. (1) using a split-operator algorithm, referred to as “multislice” in the TEM literature, as described in [29]. Here, the electrostatic scattering potentials were assembled from parameterized atomic potentials described in [39], including experimental values of the Debye-Waller factors [40, 41] and the related complex absorptive part of the potential employing a temperature of $300\,\mathrm{K}$ and an acceleration voltage of $300\,\mathrm{kV}$ [42]. The parametrized magnetic vector potentials reported in [43] were used. In order to reduce computation time, we did not simulate the thermal diffuse background due to inelastic phonon or magnon scattering in electron diffraction because that background does not affect the symmetry of elastic CBED patterns. The background is important, however, when considering whether weak magnetic diffraction effects can be reliably measured utilizing a certain electron dose in the experiment [44]. In all simulations, the starting wave function (i.e., the convergent TEM probe beam) has been normalized in intensity to unity when summed over all sampling points of the diffraction pattern. For the exemplary simulations we picked two anitferromagnetic crystals that are suited to illustrate the different aspects of the proposed method.

### v3

@@ I Introduction
Magnetic point and space groups [Zamorzaev1953], also referred to as Shubnikov groups, provide a fundamental classification underlying all magnetic or spin-related properties of crystalline solids [Rodriguez-Carvajal2019], such as ferro-, ferri- and antiferromagnetism [Blundell2001], altermagnetism [Simejkal2022], topological insulation [Tokura2019], and semimetallicity [Armitage2018, Watanabe2018]. Magnetic space groups extend the crystallographic space groups by combining spatial symmetries with time-reversal symmetry, which connects opposite spin orientations, and provide compatibility conditions for magnetic properties following Neumann’s principle. Therefore, experimental probes of magnetic point-group and space-group symmetries, analogous to probes of crystallographic symmetries, are indispensable for studies of solid-state magnetism, with neutron diffraction being the main probe [Izyumov1991]. A direct determination of the magnetic symmetry from the symmetries of a neutron diffraction pattern is, however, hampered by Friedel’s law that holds for scattering within the Born approximation: Opposite reflections in a neutron diffraction pattern are always symmetric even for noncentrosymmetric space groups. Motivated by this and other limits of neutron diffraction, such as limited spatial resolution and relatively large required crystal sizes, we establish an alternative technique based on electron diffraction in this paper. We introduce a method for the determination of magnetic point groups that allows direct measurement of symmetries with a spatial resolution reaching several nanometers, referred to as “magnetic convergent-beam electron diffraction.”
@@
While the conventional CBED technique focuses solely on structural symmetries of the crystal imprinted on the electrostatic scattering potential, a growing number of electron-scattering studies highlight the impact of the magnetic field. In an early study, Shen and Laughlin [Shen1990] showed the breaking of structural CBED symmetries by the presence of the strong uniaxial macroscopic magnetic field in ferromagnetic PrCo5. Later, Loudon [Loudon2012] revealed the existence of purely magnetic electron-diffraction spots in parallel-beam electron diffraction on antiferromagnetic NiO, where the antiferromagnetic unit cell is doubled compared to the structural one. This work paved the way for the detection of antiferromagnetic order in the Transmission Electron Microscope (TEM). Recently, Kohno et al. [Kohno2022] demonstrated the sensitivity of atomic-resolution differential phase contrast (another convergent-beam electron-diffraction technique, where first moments of diffraction patterns are analyzed) to the antiferromagnetic structure of Fe2O3. These and similar observations [Krizek2022, Nguyen2023, Tanigaki2024] suggest the experimental possibility to detect the generally weak electron-diffraction effects due to atomic magnetic fields [Edstrom2016a] in antiferromagnets, which possess no macroscopic magnetic field. The latter is the main impetus for the following generalization of the conventional CBED symmetry analysis by taking into account symmetries of the magnetic crystal structure.
@@ II Electron diffraction at electric and magnetic potentials
Electron beams accelerated by voltages in the $60\,\mathrm{kV}$ to $300\,\mathrm{kV}$ range attain relativistic velocities (e.g., $v\approx 0.8\,c$ at $300\,\mathrm{kV}$ acceleration voltage) and are predominantly scattered by very small angles (less than $100\,\mathrm{mrad}$) relative to the optical axis in a transmission electron microscope. In this small-angle, or paraxial scattering, limit, the effects of spin through spin-orbit coupling and Zeeman coupling are negligible, meaning the spinor nature of the electron wave function can be ignored. As a result, spin does not influence the diffraction, allowing us to treat the electron’s wave function as a scalar. Moreover, the resulting scalar wave function $\psi$ may be decomposed into a fast carrier wave function and an envelope wave function $\Psi$ that is slowly varying along the optical axis, i.e., $\psi(\boldsymbol{r}_{\bot},z)=e^{ik_{0}z}\,\Psi(\boldsymbol{r}_{\bot},z)$ with $\boldsymbol{r}_{\bot}=\left(x,y\right)$ denoting the in-plane slab coordinates and $k_{0}$ the wave number of the fast carrier wave. Here, the time-dependent oscillation has already been separated, i.e., we only consider eigensolutions with a certain electron energy $E$. Inserting the above decomposition into the Klein-Gordon equation for scalar relativistic particles and neglecting a couple of terms that are small due to the high beam energy, the evolution of the envelope wave function along the optical axis can be described by the paraxial Schrödinger equation [Rother2009, Edstrom2016] (see Appendix A for a detailed derivation)
@@
The symmetries of the solution of the above equation in the absence of the vector potential, $\boldsymbol{A}=0$, have been exploited for structural symmetry determination using CBED based on the following argument. A CBED pattern corresponds to the absolute square of the electron wave function in the far field behind the object (coordinates $\boldsymbol{k}_{\perp}=\left(k_{x},k_{y}\right)^{\mathrm{T}}$). Therefore, the CBED discs appear at integer multiples of unit vectors of the crystal’s reciprocal lattice according to Bragg’s law, see Fig. 1. The electron wave function, on the other hand, inherits symmetries from the paraxial Hamiltonian $\hat{H}_{\bot}$ and hence symmetries of the electric scattering potential $\Phi$, which leads to a direct relationship between symmetries of the latter and those of the CBED pattern [Bird1989]. Given the similar mathematical structure of the above equation with and without magnetic vector potential, this reasoning may be also adopted to derive the relationship between CBED pattern symmetries and magnetic point-group symmetries. Because of the additional sign change of the magnetic vector potential upon time-reversal, the latter has to be included in the symmetry considerations, when considering CBED on magnetic materials.
The structure of Eq. (1) corresponds to a two-dimensional time-dependent Schrödinger equation, i.e., a $\left(2+1\right)$-dimensional theory, with $z$ and the paraxial Hamiltonian $\hat{H}_{\bot}$ taking the place of the time coordinate and the usual Hamiltonian, respectively [Berry1971, Gratias1983]. In the following, we will take advantage of this analogy, allowing for a substantial portion of the time-dependent Schrödinger equation’s properties and solutions to be transferred to paraxial scattering.
@@
$\Psi(z)=\underbrace{Z\exp\left(-i\int_{z_{0}}^{z}dz^{\prime}\,\hat{H}_{\bot}\left(z^{\prime}\right)\right)}_{\equiv\hat{K}(z,z_{0})}\Psi(z_{0}),$ (2)
@@
In full analogy to time evolution, the $z$ evolution is unitary, conserving the norm of the wave function along $z$, because the Hamiltonian $H_{\bot}$ is hermitian. Moreover, any in-plane symmetries (rotations, mirror reflections, and translations in the $xy$ plane) leave the propagator invariant because the corresponding unitary symmetry transformations $\hat{T}_{\bot}$ commute with the Hamiltonian $\hat{H}_{\bot}$ and hence with the propagator, $\hat{T}_{\bot},\hat{K}]=0$. Translational symmetries correspond to phase factors in reciprocal space, which are not directly visible as diffraction symmetries but instead lead to absences of diffraction peaks. Hence, only in-plane rotation and mirror symmetries, i.e., the in-plane point-group symmetry of the slab, translate to symmetries of the CBED pattern, provided that the initial probe wave function also has these symmetries. Rotational and mirror symmetries are present for a circularly symmetric converged electron beam used in CBED experiments. To match these beam symmetries with the point group symmetries of the sample (i.e., to align the beam propagation axis with the high-symmetry crystallographic axis of the sample), the beam direction has to be carefully adjusted in the experiment. Here, the beam position (center of mass of the beam) within a unit cell, i.e., whether it strictly coincides with a high-symmetry axis or not, is typically irrelevant if the size of the convergent illumination spans several unit cells.
@@ III Magnetic CBED groups
Starting from the magnetic point group of the bulk crystal, we first need to construct the subgroup of elements that leave the slab geometry invariant, as discussed in Sec. II. This leads to the $10$ two-dimensional structural point groups, in Schoenflies notation, $C_{n}$ and $D_{n}$ with $n=1,2,3,4,6$. It is easy to check explicitly that all elements of these groups commute with z reversal. The fact that z reversal is antiunitary does not affect the commutation relations  [^1] One can find a purely real representation of all group elements and thereby show that the additional complex conjugation introduced by making z reversal antiunitary does not change the commutation relations.. Time reversal commutes with all possible structural transformations in any case [Wigner1931]. These arguments justify assumptions (c) and (d).
Including the antiunitary time-reversal operation $\mathcal{T}$ leads to $31$ two-dimensional magnetic point groups in total [Litvin2013]: First, the $10$ groups, $G$, noted above, which do not contain time reversal and are called “monochromatic.” Second, $10$ groups $G+G\mathcal{T}$, which result from adding time reversal $\mathcal{T}$ as a generator. These groups are called “gray.” Third, $11$ groups that are created by identifying a halving subgroup $H$ of a monochromatic group $G$ and forming the group $H+(G\setminus H)\mathcal{T}$, i.e., by multiplying the coset $G\setminus H$ by time reversal. These groups are called “dichromatic.” For an introduction to magnetic point groups see, for example, [Bradley_1968, DDJ2008].
@@
double dichromatic with $c_{n}\mathcal{T}$ $Q+Qc_{n}\mathcal{T}+Qm_{x}\Omega+Qc_{n}m_{x}\mathcal{Z}$ $\mathcal{T}$, $\mathcal{Z}$ dichromatic 33
double dichromatic with $c_{n}\mathcal{Z}$ $Q+Qc_{n}\mathcal{Z}+Qm_{x}\Omega+Qc_{n}m_{x}\mathcal{T}$ $\mathcal{T}$, $\mathcal{Z}$ dichromatic 33
double dichromatic with $c_{n}\Omega$ $Q+Qc_{n}\Omega+Qm_{x}\mathcal{T}+Qc_{n}m_{x}\mathcal{Z}$ $\mathcal{T}$, $\mathcal{Z}$ dichromatic 33
@@ IV Relation between magnetic CBED groups and magnetic point groups
We first consider a slab with surface normal oriented along the $[001]$ crystallographic axis, i.e., the fourfold symmetry axis. Thus, this slab orientation is compatible with the point group $4^{\prime}/m$. The mirror reflection in the horizontal plane is then mapped to the $z$-reversal symmetry, which we denote by $1_{R}$, following the conventions from CBED literature [Buxton1976]. The corresponding magnetic CBED group is thus $4^{\prime}1_{R}$.
On the other hand, a slab with normal along $2\bar{1}0]$ is incompatible with all operations that involve rotations by $90^{\circ}$. This reduces the symmetry of the slab to the point group $2/m$ (see Fig. 2). Describing this point group within the coordinate system defined by the slab / CBED experiment (i.e., $z$-axis parallel to beam slab normal / beam direction), the mirror symmetry along $c$ becomes a mirror along $y$ and the twofold rotation around $c$ becomes a $z$-reversal combined with a $x$-mirror. Hence, the twofold rotation is mapped to $m_{R}$. Noting that the product of $m$ and $m_{R}$ is a two-fold rotation within the plane multiplied by $z$-reversal, i.e. $2_{R}$, the magnetic CBED group of the $2\bar{1}0]$ slab turns out to be $2_{R}mm_{R}$.
These examples illustrate the general procedure we apply to generate the complete mapping of all magnetic point groups: First, we generate a representative magnetic structure for a certain magnetic point group. In the second step, we analyze the preserved symmetry operations for a certain slab orientation. Finally, we map the reversal of the coordinate normal to the slab to the antiunitary $z$-reversal. To cover all possibilities, i.e., all symmetry classes of slab orientations, we probe all classes of orientations that yield different conventional CBED symmetries, as tabulated in the literature [Buxton1976]. It turns out that all possible 125 magnetic CBED groups from Sec. III are generated by applying the above algorithm for all 122 three-dimensional magnetic point groups. The above procedure was carried out with the help of a software package provided by De Graef [DeGraef2010].
@@ V Computational examples
High-resolution TEM data as well as electron-diffraction data from magnetic crystal structures can be reliably and quantitatively reproduced by numerical integration of Eq. (1) using a split-operator algorithm, referred to as “multislice” in the TEM literature, as described in [Edstrom2016]. Here, the electrostatic scattering potentials were assembled from parameterized atomic potentials described in [Peng1996], including experimental values of the Debye-Waller factors [lee_magnetoelastic_2016, emery_variable_2011] and the related complex absorptive part of the potential employing a temperature of $300\,\mathrm{K}$ and an acceleration voltage of $300\,\mathrm{kV}$ [Castellanos-Reyes2023]. The parametrized magnetic vector potentials reported in [lyon_parameterization_2021] were used. In order to reduce computation time, we did not simulate the thermal diffuse background due to inelastic phonon or magnon scattering in electron diffraction because that background does not affect the symmetry of elastic CBED patterns. The background is important, however, when considering whether weak magnetic diffraction effects can be reliably measured utilizing a certain electron dose in the experiment [Snarski-Adamski_simulations_2023]. In all simulations, the starting wave function (i.e., the convergent TEM probe beam) has been normalized in intensity to unity when summed over all sampling points of the diffraction pattern. For the exemplary simulations we picked two anitferromagnetic crystals that are suited to illustrate the different aspects of the proposed method.